OpenStudy (anonymous):

What is the circumference of a circle with an area of 64 pi in^2? a. 8 pi in b. 16 pi in c. 1024 pi in d. 32 pi in e. none of these i got b. 16 pi in ?

OpenStudy (anonymous):

@bibby

OpenStudy (anonymous):

i also got 8 pi in but i thought 16 was correct

OpenStudy (bibby):

What is the radius of a circle with area 64 pi? well first we need the formula for area

OpenStudy (anonymous):

A=r^2

OpenStudy (bibby):

so we have \(A=\pi r^2\) \(64\cancel{\pi}=\cancel{\pi} r^2 \)

OpenStudy (anonymous):

okay and then circumference is 2 pi r

OpenStudy (bibby):

first, what is r?

OpenStudy (anonymous):

i'm getting 28.3 as my answer

OpenStudy (anonymous):

as the circumference

OpenStudy (bibby):

well not exactly

OpenStudy (bibby):

64=\(\ r^2\) \(\sqrt{64}=\ r\) r=8

OpenStudy (anonymous):

and if i plug the formula in i'm getting 50.24?

OpenStudy (bibby):

2 *pi * r = 2*8*pi = 16pi

OpenStudy (anonymous):

okay so 16 pi in is correct then?

OpenStudy (bibby):

yeah